# Appendix 0.A Quantitative Results

### v1

@@ Appendix 0.A Quantitative Results
We compute several metrics for 3D generation results with SDS[poole2022dreamfusion, wang2023score] and FSD. The hyperparameters are kept the same as those in Appendix 0.B. When using Stable Diffusion[song2020denoising] as backbone, we use 16 prompts and 4 random seeds for each prompt (Tab. 1). When using MVDream[shi2023mvdream] as backbone, we also use 16 prompts and 4 random seeds for each prompt (Tab. 2).
Table 1: Stable Diffusion[song2020denoising] as backbone
@@
Table 2: MVDream[shi2023mvdream] as backbone
@@
SDS[poole2022dreamfusion, wang2023score] $\mathbf{30.93\pm 3.40}$ $1.77\pm 0.37$ $86.6\pm 33.4$
@@ 0.A.0.1 CLIP score
We compute CLIP score[hessel2021clipscore, radford2021learning] using ViT-B/32 to measure the semantic similarity between the renderings of the generated 3D object and the input text prompt. We sample 24 views for each prompt and each seed when computing CLIP score.
@@ 0.A.0.2 IS score
We compute IS score[salimans2016improved] to measure both the image quality and diversity. We first compute the IS scores of sampled views for each prompt and then average the IS scores across prompts.
@@ 0.A.0.3 CROSS-FID score
To directly measure the diversity of generation results, it is natural to measure the inception distance between different generated samples. We first sample 24 views for each prompt and each seed. Then we separate the images corresponding to random seeds 0, 1 and 2, 3 into two sets of images. We compute FID[heusel2017gans] of the two sets of images and average the FID score across prompts. We term this score as CROSS-FID score since it is different from the standard way of using FID to evaluate GANs[heusel2017gans].
Appendix 0.B Additional Experiment Results
0.B.1 Additional Generation Results
Figure 1: Comparisons between FSD and SDS-like baseline methods. We present results of FSD and compare them to baseline methods, namely ProlificDreamer (VSD)[wang2024prolificdreamer], SDS[poole2022dreamfusion, wang2023score], CSD[ho2022classifier]. We use random seeds from 0 to 3 for each prompt.
Figure 2: Experiment results of FSD and SDS[poole2022dreamfusion, song2020score] with Stable Diffusion[song2020denoising] as backbone.
Figure 3: Experiment results of FSD and SDS[poole2022dreamfusion, song2020score] with MVDream[shi2023mvdream] as backbone.
@@ 0.B.1 Additional Generation Results
Specifically, in Fig. 1, we present results of FSD and compare them to baseline methods, namely ProlificDreamer (VSD)[wang2024prolificdreamer], SDS[poole2022dreamfusion, wang2023score], CSD[ho2022classifier]. In Fig. 2, we provide additional comparisons between SDS and FSD with Stable Diffusion[song2020denoising] as the backbone Diffusion Model. In Fig. 3, we provide additional comparisons between SDS and FSD with MVDream[shi2023mvdream] as the backbone Diffusion Model. We do not provide experiment results with Diffusion Models designed for novel view synthesis (e.g. SyncDreamer[liu2023syncdreamer]) since we found FSD does not improve the diversity of the back view of an object significantly given the front view when using Diffusion Models of such category.
0.B.2 Additional Ablations
@@
0.B.3 Additional Experiments on Noise Prior
@@
Appendix 0.C Implementation Details
0.C.1 3D Experiments using FSD
The backbone Diffusion Model for Fig. 1 in the main text is MVDream[shi2023mvdream]. The configuration for this figure is the same as Fig. 7 in the main text.
We use sqrt-annealing proposed by HiFA[zhu2023hifa] for 3D experiments in this work. We use the same CFG[ho2022classifier] scale for SDS[shi2023mvdream, wang2023score] and FSD. We use CFG=1 when applying image loss trick[zhu2023hifa] since we find the reconstructed images are unrealistic when the CFG scale is large, which is not good guidance. We follow the default setting of threestudio[threestudio2023] code base for other hyperparameters. We keep the iteration number and learning rate of FSD the same as SDS[shi2023mvdream, wang2023score] in 3D experiments. We mainly conduct our experiments using NeRF representation[muller2022instant] since SDS-like methods are not sensitive to the form of 3D representations.
0.C.2 2D Experiments using FSD
@@ Appendix 0.C Implementation Details
Methods Name SDS[poole2022dreamfusion, wang2023score] NFSD[katzir2023noise] VSD[wang2024prolificdreamer] FSD(ours) DDIM[song2020denoising]
@@
Optimizer Adam[kingma2014adam] Adam Adam Adam -
@@
Let denote the prediction of Diffusion Models as $\boldsymbol{\epsilon}_{y}^{t}=\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t}|y,t)$ and $c$ the classifier-free-guidance (CFG)[ho2022classifier] scale. Then
@@
is the loss function of SDS[poole2022dreamfusion, wang2023score], where $\emptyset$ is the empty prompt.
@@
is the loss function of NFSD[katzir2023noise], where $y_{\text{neg}}$ is a text negative prompt.
@@
is the loss function of VSD[wang2024prolificdreamer], where $\boldsymbol{\epsilon}^{t}_{\text{lora}}$ the Diffusion Model fine-tuned by LoRA[hu2021lora].
@@
Our re-implementation of $L_{NFSD}$ is slightly different from the original[katzir2023noise] design of NFSD by ignoring the condition when $t<200$, to keep the formulation simple. In this way, we find $\boldsymbol{\epsilon}^{t}_{\text{lora}}$ used in VSD[wang2024prolificdreamer] may work in a similar way as $\boldsymbol{\epsilon}^{t}_{y_{\text{neg}}}$ in NFSD[katzir2023noise]. As a result, single-particle VSD may not be able to generate diverse samples since it is still mode-seeking, aligning with the observation of ESD[wang2023taming]. But we do not conduct further investigations which are out of the scope of our work.
Appendix 0.D Theory of Flow Score Distillation
@@
0.D.1 Diffusion PF-ODE
The Diffusion PF-ODE [song2020score] can be written in the following form:
@@ 0.D.1 Diffusion PF-ODE
where $f(t)=\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t},\;g^{2}(t)=\frac{\mathrm{d}\sigma_{t}^{2}}{\mathrm{d}t}-2\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t}\sigma^{2}_{t}$, according to DPM-solver[lu2022dpm]. To get Eq. (5) in the main text, we take the derivative of $\boldsymbol{x}_{t}/\alpha_{t}$:
@@
This equation is the scaled version of Diffusion PF-ODE under the notation of Karras et al. [karras2022elucidating].
0.D.2 DDIM Sampling
We derive the DDIM[song2020denoising] sampling algorithm in this section. According to the first order discretization of Eq. 6:
@@ 0.D.2 DDIM Sampling
the sampling algorithm of DDIM[song2020denoising] can be derived as the following equation:
@@ 0.D.3 Proof of Proposition 1
[…]
@@
Appendix 0.E Visualize the Change-of-Variable
@@
(b) DDIM[song2020denoising]
Figure 7: Generation results of FSD and DDIM. We apply FSD on 2D image generation using first-order discretization Eq. 10 instead of Adam[kingma2014adam]. In this case, we find FSD is the same as DDIM[song2020denoising] except some negligible differences, which may come from the differences on handling initial conditions.
Even though we show FSD can generate similar images When using the same initial noise in the main text, there are notable differences between the generated images since FSD uses Adam[kingma2014adam] to update parameters while DDIM uses first-order discretization of Diffusion PF-ODE. We show generation results of FSD that use first-order discretization in Fig. 7.
@@
Figure 9: Visualization of $\boldsymbol{x}_{t}$, $\hat{\boldsymbol{x}}^{\text{c}}_{t}$ and $\hat{\boldsymbol{x}}^{\text{gt}}_{t}$ in FSD generation process. Compared to DDIM, FSD uses Adam[kingma2014adam] to update parameters. As a result, FSD needs more iterations to converge (See Tab. 3).
@@
Appendix 0.F Algorithm for Flow Score Distillation
@@
Appendix 0.G Practical Designing Rules for $\boldsymbol{\epsilon}(\boldsymbol{c})$
0.G.1 Practical Designing Rules
We do not specify the form of $\boldsymbol{\epsilon}(\boldsymbol{c})$ in the general form of FSD in the main text. However, it is intuitive to align $\boldsymbol{\epsilon}(\boldsymbol{c})$ in 3D space like the noise priors used in Video Diffusion Models[chang2024how, ge2023preserve, qiu2023freenoise]. We have tried several designs of $\boldsymbol{\epsilon}(\boldsymbol{c})$ and summarized several design rules for designing $\boldsymbol{\epsilon}(\boldsymbol{c})$ as well as the related potential problems if violating the design rules for $\boldsymbol{\epsilon}(\boldsymbol{c})$:
@@ 0.G.1 Practical Designing Rules
  • The noise associated with different camera views should be aligned in 3D space. Otherwise, FSD may degrade to the original SDS. This is consistent with the observation of recent work[ge2023preserve] on Video Generation, which showed that the noise maps corresponding to different video frames are highly correlated.
@@
0.G.2 World-map Noise Function

### v2

@@ Appendix 0.A Quantitative Results
We compute several metrics for 3D generation results with SDS[32, 47] and FSD. The hyperparameters are kept the same as those in Appendix 0.C. When using Stable Diffusion[43] as backbone, we use 16 prompts and 4 random seeds for each prompt (Tab. 1). When using MVDream[40] as backbone, we also use 16 prompts and 4 random seeds for each prompt (Tab. 2).
Table 1: Stable Diffusion[43] as backbone
@@
Table 2: MVDream[40] as backbone
@@
SDS[32, 47] $\mathbf{30.93\pm 3.40}$ $1.77\pm 0.37$ $86.6\pm 33.4$
@@ 0.A.0.1 CLIP score
We compute CLIP score[10, 34] using ViT-B/32 to measure the semantic similarity between the renderings of the generated 3D object and the input text prompt. We sample 24 views for each prompt and each seed when computing CLIP score.
@@ 0.A.0.2 IS score
We compute IS score[35] to measure both the image quality and diversity. We first compute the IS scores of sampled views for each prompt and then average the IS scores across prompts.
@@ 0.A.0.3 CROSS-FID score
To directly measure the diversity of generation results, it is natural to measure the inception distance between different generated samples. We first sample 24 views for each prompt and each seed. Then we separate the images corresponding to random seeds 0, 1 and 2, 3 into two sets of images. We compute FID[11] of the two sets of images and average the FID score across prompts. We term this score as CROSS-FID score since it is different from the standard way of using FID to evaluate GANs[11].
Appendix 0.B Discussions
0.B.0.1 Difference with Consistent3D
Recent work Consistent3D [52] also applied fixed noise when conducting SDS-like generation. In Consistent3D, they follow the idea of Consistent Training [44] and use the rendered image perturbed with fixed noise to approximate the starting point of the deterministic flow. In our method, for the same camera view, we also add fixed noise to the rendered image, but the noised image is used to simulate a variable in the middle of a PF-ODE trajectory, which is different from Consistent3D. Our FSD loss is also different from the CDS loss in Consistent3D, even when our view-dependent noise function gives the same noise for all camera views, implying an essential difference between our method and Consistent3D.
0.B.0.2 Connection to Signal-ODE
Our reformulated ODE (Eq. 12) is equivalent to the Signal-ODE presented in the concurrent and independent work BOOT [8], which aims to distill a fast image generator. When the diffusion model is changed to sample prediction in Eq. 12, our reformulated PF-ODE is the same as the Signal-ODE in BOOT. In BOOT, they let the student image generation model predict the clean variables $\hat{\boldsymbol{x}}^{\text{c}}_{t}$ on the ODE trajectory, while our method uses images rendered from 3D representation $\theta$ to model the clean variables $\hat{\boldsymbol{x}}^{\text{c}}_{t}$ on the ODE trajectory.
Appendix 0.C Additional Experiment Results
0.C.1 Additional Generation Results
Figure 9: Comparisons between FSD and SDS-like baseline methods. We present results of FSD and compare them to baseline methods, namely ProlificDreamer (VSD)[50], SDS[32, 47], CSD[13]. We use random seeds from 0 to 3 for each prompt.
Figure 10: Experiment results of FSD and SDS[32, 45] with Stable Diffusion[43] as backbone.
Figure 11: Experiment results of FSD and SDS[32, 45] with MVDream[40] as backbone.
@@ 0.C.1 Additional Generation Results
Specifically, in Fig. 9, we present results of FSD and compare them to baseline methods, namely ProlificDreamer (VSD)[50], SDS[32, 47], CSD[13]. In Fig. 10, we provide additional comparisons between SDS and FSD with Stable Diffusion[43] as the backbone Diffusion Model. In Fig. 11, we provide additional comparisons between SDS and FSD with MVDream[40] as the backbone Diffusion Model. We do not provide experiment results with Diffusion Models designed for novel view synthesis (e.g. SyncDreamer[24]) since we found FSD does not improve the diversity of the back view of an object significantly given the front view when using Diffusion Models of such category.
0.C.2 Additional Ablations
@@
0.C.3 Additional Experiments on Noise Prior
@@
Appendix 0.D Implementation Details
0.D.1 3D Experiments using FSD
The backbone Diffusion Model for Fig. 1 in the main text is MVDream[40]. The configuration for this figure is the same as Fig. 7 in the main text.
We use sqrt-annealing proposed by HiFA[58] for 3D experiments in this work. We use the same CFG[13] scale for SDS[40, 47] and FSD. We use CFG=1 when applying image loss trick[58] since we find the reconstructed images are unrealistic when the CFG scale is large, which is not good guidance. We follow the default setting of threestudio[9] code base for other hyperparameters. We keep the iteration number and learning rate of FSD the same as SDS[40, 47] in 3D experiments. We mainly conduct our experiments using NeRF representation[30] since SDS-like methods are not sensitive to the form of 3D representations.
0.D.2 2D Experiments using FSD
@@ Appendix 0.D Implementation Details
Methods Name SDS[32, 47] NFSD[17] VSD[50] FSD(ours) DDIM[43]
@@
Optimizer Adam[19] Adam Adam Adam -
@@
Let denote the prediction of Diffusion Models as $\boldsymbol{\epsilon}_{y}^{t}=\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t}|y,t)$ and $c$ the classifier-free-guidance (CFG)[13] scale. Then
@@
is the loss function of SDS[32, 47], where $\emptyset$ is the empty prompt.
@@
is the loss function of NFSD[17], where $y_{\text{neg}}$ is a text negative prompt.
@@
is the loss function of VSD[50], where $\boldsymbol{\epsilon}^{t}_{\text{lora}}$ the Diffusion Model fine-tuned by LoRA[14].
@@
Our re-implementation of $L_{NFSD}$ is slightly different from the original[17] design of NFSD by ignoring the condition when $t<200$, to keep the formulation simple. In this way, we find $\boldsymbol{\epsilon}^{t}_{\text{lora}}$ used in VSD[50] may work in a similar way as $\boldsymbol{\epsilon}^{t}_{y_{\text{neg}}}$ in NFSD[17]. As a result, single-particle VSD may not be able to generate diverse samples since it is still mode-seeking, aligning with the observation of ESD[48]. But we do not conduct further investigations which are out of the scope of our work.
Appendix 0.E Theory of Flow Score Distillation
@@
0.E.1 Diffusion PF-ODE
The Diffusion PF-ODE [45] can be written in the following form:
@@ 0.E.1 Diffusion PF-ODE
where $f(t)=\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t},\;g^{2}(t)=\frac{\mathrm{d}\sigma_{t}^{2}}{\mathrm{d}t}-2\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t}\sigma^{2}_{t}$, according to DPM-solver[26]. To get Eq. (5) in the main text, we take the derivative of $\boldsymbol{x}_{t}/\alpha_{t}$:
@@
This equation is the scaled version of Diffusion PF-ODE under the notation of Karras et al. [16].
0.E.2 DDIM Sampling
We derive the DDIM[43] sampling algorithm in this section. According to the first order discretization of Eq. 22:
@@ 0.E.2 DDIM Sampling
the sampling algorithm of DDIM[43] can be derived as the following equation:
@@ 0.E.3 Proof of Proposition 1
for the reverse diffusion process and then we can apply change-of-variable on $\boldsymbol{x}_{t}$. Finally
@@
Appendix 0.F Visualize the Change-of-Variable
@@
(b) DDIM[43]
Figure 15: Generation results of FSD and DDIM. We apply FSD on 2D image generation using first-order discretization Eq. 26 instead of Adam[19]. In this case, we find FSD is the same as DDIM[43] except some negligible differences, which may come from the differences on handling initial conditions.
Even though we show FSD can generate similar images When using the same initial noise in the main text, there are notable differences between the generated images since FSD uses Adam[19] to update parameters while DDIM uses first-order discretization of Diffusion PF-ODE. We show generation results of FSD that use first-order discretization in Fig. 15.
@@
Figure 17: Visualization of $\boldsymbol{x}_{t}$, $\hat{\boldsymbol{x}}^{\text{c}}_{t}$ and $\hat{\boldsymbol{x}}^{\text{gt}}_{t}$ in FSD generation process. Compared to DDIM, FSD uses Adam[19] to update parameters. As a result, FSD needs more iterations to converge (See Tab. 3).
@@
Appendix 0.G Algorithm for Flow Score Distillation
@@
Appendix 0.H Practical Designing Rules for $\boldsymbol{\epsilon}(\boldsymbol{c})$
0.H.1 Practical Designing Rules
We do not specify the form of $\boldsymbol{\epsilon}(\boldsymbol{c})$ in the general form of FSD in the main text. However, it is intuitive to align $\boldsymbol{\epsilon}(\boldsymbol{c})$ in 3D space like the noise priors used in Video Diffusion Models[3, 7, 33]. We have tried several designs of $\boldsymbol{\epsilon}(\boldsymbol{c})$ and summarized several design rules for designing $\boldsymbol{\epsilon}(\boldsymbol{c})$ as well as the related potential problems if violating the design rules for $\boldsymbol{\epsilon}(\boldsymbol{c})$:
@@ 0.H.1 Practical Designing Rules
  • The noise associated with different camera views should be aligned in 3D space. Otherwise, FSD may degrade to the original SDS. This is consistent with the observation of recent work[7] on Video Generation, which showed that the noise maps corresponding to different video frames are highly correlated.
@@
0.H.2 World-map Noise Function